A buffer solution is made using a weak acid, HA, with a pKa of 6.40. If the ratio of A– to HA is 0.010, what is the pH of the buffer?
1
Expert's answer
2019-03-28T04:53:41-0400

Ka = 10-6.4.

[A-]/[HA] = 0.01 = 10-2.

HA ⇄ H+ +A-

Ka = [H+][A-]/[HA] = 10-6.4.

[H+] = Ka/ ([A-]/[HA]) = 10-6.4/10-2 = 10-4.4.

pH = -log10[H+] = -log1010-4.4 = 4.4.